# Simulated Annealing for Quadratic and Higher-Order Unconstrained Integer Optimization

### v1

@@ 5 Summary
We proposed an SA method for optimizing polynomial functions composed of integer variables, namely the QUIO and HUIO problems. By storing and updating the coefficients required for energy-difference computations, the transition probabilities can be evaluated efficiently. We also proposed a method for estimating the initial and final temperatures from the characteristics of the model. By comparing this direct application of SA with the traditional approach of converting the problems into QUBO or HUBO formulations, we showed that the direct approach achieves better performance in practice.
@@
When the bounds of the integer variables are wide, the Metropolis method, which selects the next state randomly, has difficulty finding the optimal solution, while the heat bath method becomes time-consuming because it must evaluate all possible states of each variable. To address these issues, we introduced the optimal-transition Metropolis method. Comparing these transition probabilities, we found that all methods tend to reach the optimal solution when the annealing time is sufficiently long, but the optimal-transition Metropolis method shows a clear advantage over the others in the short-annealing-time regime. This is particularly important in real-world applications, where the annealing time is often limited and the ability to obtain reasonably good solutions quickly is essential. At the same time, this method does not lose effectiveness in the long-time regime, converging to the same solutions as the other transition probabilities.
@@
In this study, the optimal-transition Metropolis method was implemented only in cases where each variable appears with exponent at most four in every term. However, most practical optimization problems involve objective functions and constraints that are at most quadratic. Even when quadratic constraints are embedded into the objective function through penalty or augmented Lagrangian terms[Bertsekas1982], the resulting expressions still contain only up to fourth-order terms, so this restriction is sufficiently practical for real-world applications.
@@
Based on these observations, the choice of transition probability can be summarized as follows. When the objective function consists only of multilinear terms, the heat bath method can be executed efficiently using the inverse-transform approach, so either the heat bath method or the optimal-transition Metropolis method is appropriate. For general polynomial objectives, the optimal-transition Metropolis method is preferable if it can be used. Otherwise, the conventional Metropolis method becomes the practical choice.
@@
Our results obtained in this study suggest that the proposed approach provides a new and efficient solution method for general integer polynomial optimization problems, into which many real-world tasks can naturally be formulated. Moreover, the optimal-transition mechanism introduced in this work can be applied without modification to SA for continuous variables, and the underlying idea may also prove effective for metaheuristics beyond SA.
@@
Finally, we highlight several directions for future work. Concerning transition probabilities, other efficient update schemes[aoap/1177005652, biomet/83.3.681, PhysRevE.70.056705, PhysRevLett.105.120603, 10.1137/110832288] are also well known. As in the case of the heat bath method, it is essential to compute transition probabilities efficiently when the variable bounds are large. If these methods are incorporated appropriately into SA, they may further enhance its performance. Regarding update strategies, this work focused on single-variable updates, where one variable is updated at a time. Incorporating cluster-based updates, such as those used in the Swendsen–Wang[PhysRevLett.58.86] or Wolff[PhysRevLett.62.361] algorithms, or allowing simultaneous updates of multiple variables, may offer additional performance improvements.

### v2

@@ 5 Summary
We proposed an SA method for optimizing polynomial functions composed of integer variables, namely the QUIO and HUIO problems. By storing and updating the coefficients required for energy-difference computations, we can evaluate the transition probabilities efficiently. We also proposed a method for estimating the initial and final temperatures from the characteristics of the model. By comparing this direct application of SA with the traditional approach of converting the problems into QUBO or HUBO formulations, we showed that the direct approach achieves better performance in practice.
@@
When the range of the integer variables are wide, the Metropolis method, which selects the next state randomly, has difficulty finding the optimal solution, while the heat bath method becomes time-consuming because it must evaluate all possible states of each variable. To address these issues, we introduced the optimal-transition Metropolis method. Comparing these transition probabilities, we found that all methods tend to reach the optimal solution when the annealing time is sufficiently long, but the optimal-transition Metropolis method shows a clear advantage over the others in the short-annealing-time regime. This is particularly important in real-world applications, where the annealing time is often limited and the ability to obtain reasonably good solutions quickly is essential. At the same time, this method does not lose effectiveness in the long-time regime, converging to the same solutions as the other transition probabilities.
@@
In this study, the optimal-transition Metropolis method was implemented only in cases where each variable appears with exponent at most four in each term. However, most practical optimization problems involve objective functions and constraints that are at most quadratic. Even when quadratic constraints are embedded into the objective function through penalty or augmented Lagrangian terms[Bertsekas1982], the resulting expressions still contain only up to fourth-order terms, so this restriction is sufficiently practical for real-world applications.
@@
Based on these observations, the choice of transition probability can be summarized as follows. When the objective function consists only of multilinear terms, the heat bath method can be executed efficiently using the inverse-transform approach, and the optimal-transition Metropolis method can also be applied without difficulty. Therefore, either the heat-bath method or the optimal-transition Metropolis method is appropriate in this case. For general polynomial objectives, the optimal-transition Metropolis method is preferable for the cases implemented in this study, where each variable appears with exponent at most four. Otherwise, the conventional Metropolis method is generally the practical choice.
@@
Our results obtained in this study suggest that the proposed approach provides a new and efficient solution method for general integer polynomial optimization problems, which many real-world tasks can naturally be formulated into. Moreover, the optimal-transition mechanism introduced in this work can be applied without modification to SA for continuous variables, and the underlying idea may also be effective for metaheuristics beyond SA.
@@
Finally, we highlight several directions for future work. Concerning transition probabilities, other efficient update schemes[aoap/1177005652, biomet/83.3.681, PhysRevE.70.056705, PhysRevLett.105.120603, 10.1137/110832288] are also well known. When the variable range is wide, it is essential for any transition-probability scheme to compute probabilities efficiently. If these methods are incorporated appropriately into SA, they may further enhance its performance. Regarding update strategies, our work focused on single-variable updates, where one variable is updated at a time. Incorporating cluster-based updates, such as those used in the Swendsen–Wang[PhysRevLett.58.86] or Wolff[PhysRevLett.62.361] algorithms, or allowing simultaneous updates of multiple variables, may offer additional performance improvements.